OpenStudy (anonymous):

what is the exact, fractional value of sin75

OpenStudy (anonymous):

sin (30+45)

OpenStudy (anonymous):

so then use sum/difference formula

OpenStudy (anonymous):

what is the sum/difference formula?

myininaya (myininaya):

gj light! thats the hardest past coming up with two numbers that =75 that have exact answers

OpenStudy (anonymous):

so sin30cos45+cos30sin45

OpenStudy (anonymous):

http://www.mathwords.com/s/sum_difference_identities.htm to strive5

myininaya (myininaya):

sin(x+y)=sinxcosy+sinycosx sin(x-y)=sinxcosy-sinycosy cos(x+y)=cosxcosy-sinxsiny cos(x-y)=cosxcosy+sinxsiny

OpenStudy (anonymous):

take it from there

OpenStudy (anonymous):

thank you that was just on the A2 Trig regents i took today and i had no idea how to do it

OpenStudy (anonymous):

doesnt sound like A2 to me more like actual trig

OpenStudy (anonymous):

there were a lot of weird questions on this regents. it was a lot harder than the past regents that i took as practice.

OpenStudy (anonymous):

fyi the answer is (sqrt(2)+sqrt(6))/4

OpenStudy (anonymous):

i think i put sqrt(2)/4 just from trying to do it with no idea of the actual formula. i might get some credit. it was a 4 point question